17-20
Answer Question 17 (1 point In order to simulate the demand for year 2020, you generate a random number of 0.445. What would be the corresponding Number of Sales? of Sales Frequency(Days)
Your Answer Question 17 options: Answer Question 18 (1 point) A restaurant sells pizza at a rate of $15/slice. Expenses for the restaurant include raw material for pizza at $9 per slice, $186 as monthly rental and $56.00 monthly as insurance. How many slices should the restaurant sell in a month to break even? Answer rounded to 0 decimal points following standard rules for rounding Your answer: Question 18 options: Answer Question 19 (1 point) Consider the following payoff table: DA: Decision Alternative Decision Alternative State of Nature ood Bad babilities 0.6 0. What is the Expected value of perfect Information (EVPI)? (Please keep 1 decimal for your answer) Your Answer: Question 19 options Answer Question 20 (1 point) All cars have at least two doors, BMW is a car hence it has at least two doors. What type of reasoning is this?
Answer #1

17. The CDF closest to the random number 0.445 is 0.319

The corresponding number of sales >=450

The correct answer is 450

E4 Number of sales Frequency Probability Cummulative probabilit 100 14 12 20 0.14 0.20 0.26 0.29 0.14 0.06 0.06 0.03 0.03 4 3

18. Break even point =Fixed costs/Price-Variable costs

=Monthly rental+Insurance /(15-9)

=186+56/(6)

=242/6

=40.33

=41 (rounding to next whole number)

19.

Expected value with perfect information EVwPI =0.6*5+0.4*2

=$ 3.8

EVwoPI=highest EMV from the decision tree =EMV of alternative 2=3.2

(we take highest because it is the profit problem)

EVPI=absolute(EVwPI-EVwoPI)=absolute(3.8-3.2)=$ 0.6

IF(ABS(1-SUM(N18,N23))0.00001,SUM(N18*21,N23*O26), NA()) K23 L M 0.6 Good 10 DA 1 12 13 14 15 16 17 0.4 Bad 3.2 0.6 Good 4 21

20. This is an example of deductive reasoning.